OpenStudy (ledah):

Solving systems of equations -x+y=12 x+2y=3

OpenStudy (campbell_st):

addthe 2 lines together to eliminate x you get 3y = 15 then y = 5 substitiute it into the 1st equation - x + 5 = 12 - x = 7 then x = -7